Given the following systems of equations, which statement is true?
y=3x-8
5x+y=19
Choices:
A) 3x-8=5x+19
B) 3x-8=19-5x
C) 5x-19=3x-8
D) 5x+3x=11

Answer by pwac(253)

Substitute the given value of y into the equation 5x + y = 19
This gives:
5x + 3x-8 = 19 Subtract 5x from both sides
which leaves:
3x-8 = 19-5x
which corresponds with statement B)
Peter